Please show how to get the answer as well!

On January 2, 2017, Oriole Company began construction of a new citrus processing plant. The automated plant was finished and ready for use on September 30, 2018. Expenditures for the construction were as follows: January 2, 2017 September 1, 2017 December 31, 2017 March 31, 2018 September 30, 2018 603000 1810800 1810800 1810800 1218000 Oriole Company borrowed $3300000 on a construction loan at 10% nterest on January 2. 2017. This loan was outstanding during the construction period e c mpany als nad nss 0 % bonds outstanding in 2017 and 2018. The interest capitalized for 2017 was: $158200 $482340 $ 474600 $120660
Answer #1
Actual interest expenditure incurred:
Principal Interest % Interest
expenditure
1 2 1*2
Loan at 10%-specific 3300000 10% 330000
7% bond-General 13560000 7% 949200
1279200
Avoidable interest=Weighted average accumulated expenditure*interest rate=3000000*9%=270000
Weighted average accumulated expenditure:
Date Expense Capitalization
period
Weight Weighted
expenditure
A B C=B/12 A*C
2-Jan 603000 12 months 1.00 603000
1-Sep 1810800 4 months 0.33 603600
31-Dec 1810800 0 months 0.00 0
1206600
Weighted average interest rate for general purpose notes=7%
Here,weighted average expeenditure is much less than the principal amount of specific loan and general purpose note.So,entire weighted expenditure is applied towards specific loan.
Avoidable interest:
Principal Interest % Interest
expenditure
Loan at 10%-specific 1206600 10% 120660
Interest to be capitalized=Lower of actual interest or avoidable interest=Lower of 1279200 or 120660=120660
Answer is $ 120660